OpenStudy (anonymous):

Seventeen times n divided by 153 equals 5. What is 75% of n?

OpenStudy (whpalmer4):

17 * n / 153 = 5. Multiply both sides by 153 to get 17n = 153*5 = 765. Divide both sides by 17 to get n = 45. 75% of n = 75/100*45 = 33.75.